Consider a population of 6 units with values 1, 2, 3, 4, 5 and 6. Write down all possible samples of size 2 from this population. Obtain the sampling distribution of the sample mean. Compute mean and variance of the sampling distribution obtained.
1
Expert's answer
2020-04-13T09:18:00-0400

We randomly choose two unts, without replacement. The order doesn't matter.


"\begin{matrix}
 Sample & sample\ mean, \ \bar{x} \\
 1,2 & 1.5 \\
 1,3 & 2\\
 1,4 & 2.5 \\
 1,5 & 3 \\
 1,6 & 3.5 \\
 2,3 & 2.5 \\
 2,4 & 3 \\
 2,5 & 3.5 \\
 2,6 & 4\\
 3,4 & 3.5 \\
 3,5 & 4 \\
 3,6 & 4.5 \\
 4,5 & 4.5 \\
 4,6 & 5 \\
 5,6 & 5.5
\end{matrix}"


"\begin{matrix}
 \bar{x} & f & probability \\
 1.5 & 1 & 1/15 \\
 2.0 & 1 & 1/15\\
 2.5 & 2 & 2/15\\
 3.0 & 2 & 2/15 \\
 3.5 & 3 & 1/5\\
 4.0 & 2 & 2/15\\
 4.5 & 2 & 2/15\\
 5.0 & 1 & 1/15\\
 5.5 & 1 & 1/15\\
 
\end{matrix}"


"\mu_{\bar{x}}=1.5({1\over15})+2.0({1\over15})+2.5({2\over15})+3.0({2\over15})+3.5({1\over5})+"

"+4.0({2\over15})+4.5({2\over15})+5.0({1\over15})+5.5({1\over5})=3.5"


"(1.5)^2({1\over15})+(2.0)^2({1\over15})+(2.5)^2({2\over15})+(3.0)^2({2\over15})+(3.5)^2({1\over5})+"

"+(4.0)^2({2\over15})+(4.5)^2({2\over15})+(5.0)^2({1\over15})+(5.5)^2({1\over5})="

"={161\over 12}"

"Var(\bar{x})=\sigma_{\bar{x}}^2={161\over 12}-(3.5)^2={7\over 6}\approx1.166667"